y = + x2 + x-1 and x is a positive number.

a Write y as a mixed number when

i x = 1 ii x = 2 iii x = 3


1
Expert's answer
2021-08-09T13:13:40-0400
"y=x^2+x^{-1}"

a.

i. "x=1"


"y=1^2+1^{-1}=2"

ii. "x=2"


"y=2^2+2^{-1}=4\dfrac{1}{2}"

iii. "x=3"


"y=3^2+3^{-1}=9\dfrac{1}{3}"